# Autonomous Learning with High-Dimensional Computing Architecture Similar to von Neumann’s

### v1

@@ 4.2 Short-Term Working Memory
There is more to consider when computing with vectors in superposition: just how many vectors to encode into a single vector? The number is limited by vector dimension and determines the amount of structure that a single vector can hold (Gallant & Okaywe 2013, Frady et al. 2018), beyond which it is necessary to store vectors in the long-term memory. Psychologists call it chunking.
The human mind can recall large amounts of data that has been organized in chunks (and in chunks of chunks, and so on). George Miller’s (1956) “Magical Number Seven, Plus or Minus Two” is the classic on the topic. The working memory combines pieces of data into chunks and stores them in the associative memory, and recalls chunks into the working memory, to compose them into further chunks or to break them down into their constituent parts.
Miller’s magical number suggests that a chunk should combine fewer than ten items. That is also about the number of vectors that can be encoded into a single vector and readily decoded with the vector math. With psychology and the vector math as clues, we assume that a working memory can hold and operate on up to about ten vectors at a time.
@@ 5 Modeling the World with Vectors
To help us understand intelligence in animals (and in ourselves), we model them with robots whose behavior is controlled by “brains” that compute with high-dimensional vectors. A robot has sensors and actuators for interacting with the world, an associative memory for storing a model of the world, and a working memory where the world model is constructed and interpreted with the vector math (Teeters et al. 2023).
@@ 5.1 The Focus
The contents of the working memory are summarized in a single $H$-dimensional vector called the focus. The sensors feed into the focus, the activators are driven from the focus, the memory is addressed by the focus, and vectors in and out of the memory pass through the focus. The vector in the focus is the system’s (robot’s) state and it represents the robot’s subjective experience at the moment—the focus is the robot’s subjective self.
@@
The focus is a theoretical construct meant to capture the idea of an integrated self. For example, any of a number of things can bring to our minds a specific experience from the past. It could be a specific sight or sound or smell or ache or pain, or some combination of them. Thus any one of them can serve as an address to memory, to evoke the past experience. Similarly, we recognize a person by how they look or sound or move or act: all those different modalities point to the same person.
@@
It is befitting that the focus is limited to a single vector because our minds can attend to only a few things at a time (cf. Miller’s magical number 7). Limited capacity is behaviorally realistic. Furthermore, because the focus is at the crossroads of sensors, actuators, and memory, an experience created by the world through the senses can also be created from the memory. In computing terms, they reside in a common $H$-dimensional space, allowing the history to be stored as a pointer chain.
@@
As an added benefit, representing sensory modalities and action in common mathematical space allows generic algorithms to be used for learning. It also means that we need to turn physical (and chemical) signals with vastly different properties (sight, sound, taste, smell, temperature, pressure, etc.) into $H$-dimensional vectors for the working memory. Best ways to do it will be worked out over time. Meanwhile, standard signal-processing techniques together with random projections have been used successfully. For example, letters of the alphabet have been represented by random vectors for written-language identification (Joshi et al. 2017), words of a language have been represented by random vectors for comparing word meaning (Kanerva et al. 2000), and Fourier coefficients for speech have been randomly projected to the $H$-dimensional space for speaker identification (Huang et al. 2022).
@@
Looking ahead, we need to learn how nervous systems and brains prepare sensory input for further processing, and build that into artificial systems. For example, the cochlea of the inner ear analyses sound into frequencies—it Fourier-transforms the sound before passing it on to the rest of the brain. Regarding vision, the optic nerve brings in information along about 1.4 million fibers and the primary visual cortex distributes it among 280 million neurons—a 200-fold increase (the numbers are from Wikipedia).
@@ 5.2 The Passage of Time
Short sequences, up to ten or so, can be encoded into a single vector in the working memory and treated as a chunk. Joshi et al. (2017) encoded short sequences of letters ($N$-grams) as chunks and used them to identify languages. The long-term associative memory makes it possible to deal with longer sequences by storing the history as a life-long pointer chain (see Traditional Computing with Numbers above). The vector for the present moment—the present state, the focus—can then be used as an address to trigger the recall of similar past episodes.
@@ 5.3 Sensing and Acting
Prediction can be accomplished by storing the system’s history in the memory as a pointer chain, as discussed above. When a vector for a moment recurs and is used to address the memory, it recalls—and predicts—the next moment. A moment later the prediction can be compared to what actually happens—this is essentially anomaly detection that can initiate further action.
@@
What is the individual’s relation to its model of the world: how is the model used? We can think of the model as a database that the individual consults, to evaluate alternative actions and to choose one that is appropriate for the present situation. Most of today’s robots work like that, but it is not what actually happens in brains because it assumes a high-level executive (a homuncules) that runs the show. The use of the model by the individual needs to be more organic. Looked at in that way, rather than an individual having a model of the world, the individual is the model.
Interaction with the world and its individuals are the data for the world model. An individual’s interaction is experienced as a sequence of states of the focus that include both sensory and motor components. When stored in memory as a pointer chain, it allows the individual to predict and to act according to the prediction. The details are worked out in the target article (Kanerva 1988a) and will not be repeated here. We merely comment on it below.
6 The Organization of an
Autonomous Learning System (1988)
@@ 6 The Organization of an
Computing with high-dimensional vectors tries to capture the organization of animals that learn and adapt. It is based on the idea that an associative memory can learn predictive modeling of the world with a circuit that is strikingly similar to the cerebellum’s. In psychologists’ terms, the associative memory is a long-term data store.
@@
The 1988 paper has nothing concrete to suggest about how the working memory actually works; it merely refers to an “Encoding Problem” as yet to be solved. The gist of the encoding problem was solved in the 1990s by computing with vectors in superposition. Other than that, the architecture is conventional (von Neumann) and suitable for both symbolic (Gayler 2003) and probabilistic (Joshi et al. 2017) computing.
The 1988 paper refers to varying the amount of information that a given data source or destination contributes to the focus vector, but does not tell us how actually to do it. Superposition suggests an answer. When each data source and destination is encoded into an $H$-dimensional vector, the vectors can be weighted by importance before adding them to the focus. For example, we can prioritize sight over sound, or vice versa, depending on the context. And so we would no longer refer to feature vectors, as that suggests vectors being partitioned into fields for features. Every “feature” that is encoded into a vector is present in every coordinate. In fact, what we traditionally think of as features are more like afterthoughts, to help us compartmentalize the world. If everything in the world were green, color as a feature would not exist.
7 Where to Now?
Computing with high-dimensional vector in superposition seems particularly suited for simulating—and possibly understanding— intelligence of animals with advanced sensory–motor systems and brains. Much needs to be worked out, however, to develop the idea into a viable technology. Fortunately, developing and testing does not require excessive hardware or computing. Ample opportunity to experiment, innovate and invent is within reach. Here are some of the challenges we face.
Programming and Control. Whether and how to do the actual programming with superposed vectors is unclear. For now, conventional programs running in conventional computers control the operation of the hyperdimensional computer in simulation or prototype.
@@ 7 Where to Now?
Associative Memory. Computing with vectors depends crucially on a high-capacity associative memory for high-dimensional vectors (a high-D RAM), that does a fast near-neighbor search in a huge and sparsely populated address space. When vectors are stored in an associative memory with themselves a addresses (called autoassociative) they can become fix points with basins of attraction of a dynamical system, resulting in rapid retrieval of stored vectors. The sparse distributed memory (Kanerva 1988b) demonstrates the concept but is inefficient. The circuits of the cerebellum can suggest more efficient designs, for example by judicious use of randomness and sparseness.
Noise. One of nature’s marvels is how we recognize a person or an object or a place or a situation—anything—as being the same now as at some earlier times, and distinct from other things, in spite of the data arriving at the senses never repeating exactly. The tendency to categorize things based on what we expect of them, seems to be built in. But even “built in” needs explaining: a physical process that sustains the behavior. Without going into it here, we merely presume that high dimensionality plays an important role, and instead look at how variability in input is handled in cognitive modeling today.
@@
Another source of “noise” cannot be ignored, and it has to do with encoding and decoding of superposed vectors. Multiplication and permutation (of binary vectors) are invertible and preserve information, but addition is only approximately invertible, and so the decoding of a vector sum puts out noisy vectors. That sets a limit on the number of vectors to include in a sum, and the need to clean up vectors decoded from a sum before further operating on them.
Neuroscience. We have a reasonable idea of how a long-term memory in the brain is organized (cf. the cerebellum) but none for the working memory. If computing with vectors produces human-like or animal-like behavior with energy efficiency comparable to the brain’s, we can hope to understand and identify brain’s circuits for working memory.
Psychology and Psychophysics. Psychologists have probed human (and animal) performance in countless experiments, and conjectured about processes that underlie memory recall, response time, attention span, interference, ability to discriminate, and so on. Relating their observations to computing with high-dimensional vectors can deepen our understanding of the underlying processes and lead to the engineering of more lifelike robots, for example.
Modeling of brains with high-dimensional vectors also opens new vistas in experimental psychology. As the theory and models improve, they will suggest experiments to confirm or counter hypothesized underlying mechanisms. This can benefit both psychology and engineering.
Language and logic are the apex of symbolic behavior and have even been considered as defining intelligence. Symbolic processing also formed the early core of AI—until we tried to build systems that learn by interacting with the world. Where logic and language failed, artificial neural nets were thought to succeed, leading to today’s deep neural nets. Impressive as they are, their energy requirements so exceed the efficiency of brains, as to rule them out as models of how brains work and make us intelligent.
@@ 8 Summary
The paper outlines a traditional (von Neumann-like) architecture for computing with high-dimensional vectors and argues for its use for autonomous learning in real time by systems such as robots. It is also meant to help us understand how humans and animals learn. It combines ideas from psychology (short-term working memory and long-term data store, chunking), neuroscience (cerebellum-like circuit for an associative memory), computer science and engineering (central processing unit, random-access memory, sequence as a pointer chain), and mathematics (algebra of vector operations, effective orthogonality of random vectors). Peculiar about computing with vectors is that it takes place in superposition. Variables and values are vectors of a common mathematical space, and so is a variable having a certain value. The memory for storing the vectors is also addressed by the vectors—it’s a “high-dimensional RAM.”
The subjective state of the system is defined by a single vector called the focus. It summarizes the contents of the working memory. The sensors feed into the focus, the actuators are driven from the focus, and the memory is accessed through the focus. The passage of time is represented by a sequence of vectors in the focus. Stored in the memory as a pointer chain, it becomes a predictive model of the world.
@@
The architecture is shallow. The memory has one fixed layer for activating memory locations and one variable layer for storing and recalling data in the activated locations. The fixed layer corresponds to the address-decoder circuit of a computer memory (RAM), and the variable layer corresponds to memory stacks that store the data. This architecture is suggested by the wiring of the cortex of the cerebellum, which contains over half the brain’s neurons and synapses. The variable layer can be trained in real time by the perceptron learning rule, and learning can happen in only a few rehearsals.
The vector math is carried out in the working memory (analogue of the Central Processing Unit, CPU) with an instruction set consisting of coordinatewise addition and multiplication, and the permutation of coordinates. Similarity of vectors is based on the dot product. The operations can combine symbolic processing and probabilistic learning.
Interaction with the world happens through sensors and actuators. Input is made into high-dimensional vectors for the working memory/focus with sensor-specific pre-processing, and the focus is decoded for commands to actuators with motor-specific post-processing. Pre- and post-processing can be engineered, inherited or learned early in the system’s existence in multiple passes with backpropagation.
@@
Conflict of Interest Statement. The author declares that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions. PK is the sole author of the article.
Funding. This work is supported in part by CoCoSys, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.
Acknowledgments. Special thanks to professor Bruno Olshausen and his Redwood Center for Theoretical Neuroscience, for providing a fertile setting for the development of ideas expressed in this paper.
Publisher’s Note. All claims expressed in this article are solely those of the authors and do not necessarily represent those of their affiliated organizations, or those of the publisher, the editors and the reviewers. Any product that may be evaluated in this article, or claim that may be made by its manufacturer, is not guaranteed or endorsed by the publisher.
Supplemental Material. RIACS Report 88.14 on the Organization of an Autonomous Learning System (Kanerva 1988a), referred to here as the Target Article.

### v2

@@ 5.2 Short-Term Working Memory
There is more to consider when computing with vectors in superposition: just how many vectors to encode into a single vector? The number is limited by vector dimension and determines the amount of structure that a single vector can hold (Gallant & Okaywe 2013; Frady et al. 2018), beyond which it is necessary to store vectors in the long-term memory. Psychologists call it chunking.
The human mind can recall large amounts of data that has been organized in chunks (and in chunks of chunks, and so on) and stored in memory. George Miller’s “Magical Number Seven, Plus or Minus Two” is the classic on the topic (Miller 1956). The working memory combines pieces of data into chunks and stores them in the associative memory, and recalls chunks into the working memory, to compose them into further chunks or to break them down into their constituent parts.
Miller’s magical number suggests that a chunk should combine fewer than ten items. That is also about the number of vectors that can be encoded into a single vector and readily decoded with the vector math. The number grows nearly linearly with vector dimensionality $D$ (Frady et al. 2018). With psychology and the vector math as clues, we assume that a 10,000-dimensional working memory can hold and operate on up to ten vectors at a time.
@@ 6 Modeling the World with Vectors
To help us understand intelligence of animals (and of ourselves), we liken them (and us) to robots whose behavior is controlled by “brains” that compute with wide vectors. Albus (1981) makes a case for humanlike robots, and Kanerva (1988b), Plate (1991) and Piantadosi et al. (2024) argue for representing concepts with high-dimensional vectors. The robot has sensors and actuators for interacting with the world, an associative memory for storing a model of the world, and a working memory where the world model is constructed and interpreted with the vector math (Teeters et al. 2023).
@@ 6.1 The Focus
The contents of the working memory are summarized in a single $D$-dimensional vector called the focus. The idea of a focus at the core of a cognitive computing architecture goes back to 1988 (Kanerva 1988a, b) but has not been adopted widely. However, it deserves a new look now that von Neumann-like computing with vectors in superposition has become understood Plate (1991). The sensors feed into the focus, the activators are driven from the focus, the memory is addressed by the focus, and vectors in and out of the memory pass through the focus. The vector in the focus is the system’s (robot’s) state and it represents the robot’s subjective experience at the moment—the focus is the robot’s subjective self (an objective state would have to include the contents of memory as well).
@@
The focus is a theoretical construct, meant to capture the idea of an integrated self. For example, any of a number of things can bring to mind a specific experience from the past; the trigger can be a specific sight or sound or smell or ache or pain, or some combination of them. Thus any one of them can serve as an address to memory, to evoke the past experience. Similarly, we recognize a person by how they look or sound or move or act: all those different modalities refer to—and retrieve from memory—the same person.
@@
It seems befitting that the focus is limited to a single vector because our minds can attend to only a few things at a time (cf. Miller’s magical number 7). Limited capacity is behaviorally realistic. Furthermore, because the focus is at the crossroads of sensors, actuators, and memory, an experience created by the world through the senses can also be created from the memory. In computing terms, they reside in a common $D$-dimensional space, allowing the history to be stored as a linked list; see subsection 3.2.
@@
As an added benefit, representing sensory modalities and action in common mathematical space allows generic algorithms to be used for learning. It also means that we need to turn physical (and chemical) signals with very different properties (sight, sound, taste, smell, temperature, pressure, etc.) into $D$-dimensional vectors for the working memory. The best ways to do it will be worked out over time. Meanwhile, standard signal-processing techniques together with random projections have been used successfully. For example, letters of the alphabet have been represented by random vectors for written-language identification (Joshi et al. 2017), words of a language have been represented by random vectors for comparing word meaning (Kanerva et al. 2000), and Fourier coefficients for speech have been randomly projected to a $D$-dimensional space for speaker identification (Huang et al. 2022).
@@
Looking ahead, we need to learn how nervous systems and brains prepare sensory input for further processing, and build that into artificial systems. For example, the cochlea of the inner ear analyses sound into frequencies—it Fourier-transforms the sound before passing it on to the rest of the brain. Regarding vision, the optic nerve brings in information along about 1.2 million axons per hemisphere (Sterling & Demb 2004) and the primary visual cortex distributes it among some 140 million neurons (Leuba & Kraftsik 1994)—a 100-fold increase.
@@ 6.2 The Passage of Time
Short sequences, up to ten or so, can be encoded into a single vector in the working memory and treated as a chunk. Joshi et al. (2017) encoded short sequences of letters ($N$-grams) as chunks and used them to identify languages. The long-term associative memory makes it possible to deal with longer sequences by storing the history in the long-term memory as a life-long linked list; see subsection 3.2. The vector for the present moment—the present state, the focus—can then be used as an address to trigger the recall of similar past episodes.
@@ 6.3 Sensing and Acting
Prediction can be accomplished by storing the system’s history in the memory as a linked list, as described in subsection 3.2. When a vector for a moment recurs and is used to address the memory, it recalls—and predicts—the next moment. A moment later the prediction can be compared to what actually happens—this is essentially anomaly detection that can initiate further action.
@@
What is the individual’s relation to its model of the world: how is the model used? We can think of the model as a database that the individual consults, to evaluate alternative actions and to choose one that is appropriate for the present situation. Most of today’s robots work like that, but it is not what actually happens in brains because it assumes a high-level executive (a homunculus) that runs the show. The use of the model by the individual needs to be more organic. Looked at in that way, rather than an individual having a model of the world, the individual is the model (and the model is the individual).
Interaction with the world and its individuals are the data for the world model. An individual’s interaction is experienced as a sequence of states of the focus that include both sensory and motor components. When stored in memory as a linked list, it allows the individual to predict and to act according to the prediction. The details are worked out in the target article (Kanerva 1988a) and will not be repeated here. We merely comment on it.
7 Autonomous Learning (Kanerva 1988a)
@@
Computing with wide vectors tries to capture the organization of animals that learn and adapt. It is based on the idea that an associative memory can learn predictive modeling of the world with a circuit that is strikingly similar to the cerebellum’s. In psychologists’ terms, the associative memory is a long-term data store.
@@
The 1988 paper has nothing concrete to suggest about how the working memory actually works. The paper merely refers to an “Encoding Problem” as yet to be solved. The gist of the encoding problem was solved by Plate by computing in superposition with Holographic Reduced Representation Plate (1991). Other than that, the architecture is conventional (von Neumann) and suitable for both symbolic (Gayler 2003) and statistical computing (Joshi et al. 2017).
The 1988 paper refers to varying the amount of information that a given data source or destination contributes to the focus vector, but does not tell us how actually to do it. Superposition suggests an answer. When each data source and destination is encoded into a $D$-dimensional vector, the vectors can be weighted by importance before adding them to the focus. For example, we can prioritize sight over sound, or vice versa, depending on the context. We would no longer refer to feature vectors, as that suggests vectors being partitioned into fields for features. Every “feature” that is encoded into a vector is present in every coordinate. In fact, what we traditionally think of as features are more like afterthoughts, to help us compartmentalize the world. If everything in the world were green, color as a feature would not exist.
8 What Next?
Computing with wide vectors seems particularly suited for simulating—and possibly understanding—intelligence of animals with advanced sensorimotor systems and brains (Albus 1981; Kanerva 1988b; Piantadosi et al. 2024). Much needs to be worked out, however, to develop the idea into a viable technology. Fortunately, developing and testing does not require excessive hardware or computing. Ample opportunity to experiment, innovate and invent is within reach (see Yang 2023, for an example). Here are some of the challenges we face.
Programming and Control. Whether and how to do the actual programming with superposed vectors is unclear. For now, conventional programs running in conventional computers control the operation of the hyperdimensional computer in simulation or prototype. However, people are beginning to think also about vector-based control structures (Tomkins-Flanagan & Kelly 2024, e.g.).
@@ 8 What Next?
Associative Memory. Computing with vectors depends crucially on a high-capacity associative memory for wide vectors (a high-dimensional RAM), that does a fast near-neighbor search in a huge and sparsely populated address space. When vectors are stored in an associative memory with themselves a addresses (called autoassociative) they can become fixed points with basins of attraction of a dynamical system, allowing rapid storing and retrieving of vectors. The Sparse Distributed Memory (Kanerva 1988b) demonstrates the concept but is inefficient. The circuits of the cerebellum can suggest more efficient designs, for example by judicious use of randomness and sparseness (Jaeckel 1989).
Noise. One of nature’s marvels is how we recognize a person or an object or a place or a situation—anything—as being the same now as at some earlier time, and distinct from other things, in spite of the data arriving at the senses never repeating exactly. The tendency to categorize things based on what we expect of them, seems to be built in. But even “built in” needs explaining: a physical process that sustains the behavior. Without going into it here, we merely presume that high dimensionality plays an important role, and instead comment on how variability in input is handled in cognitive modeling today.
@@
Another source of “noise” cannot be ignored. It has to do with encoding and decoding of superposed vectors. For binary vectors, multiplication and permutation are invertible and preserve information, but addition is only approximately invertible, and so the decoding of a vector sum puts out noisy vectors. That limits the number of vectors included in a sum, and calls for cleaning up of vectors decoded from a sum before further operating on them.
Neuroscience. We have a reasonable idea of how a long-term memory in the brain is organized (Albus 1971; Kanerva 1993; Llinás et al. 2004; Jaeckel 1989, cf. the cerebellum:) but none for the working memory. If computing with vectors produces human-like or animal-like behavior with energy comparable to brain’s, we can hope to understand and identify brain’s circuits for working memory.
Psychology and Psychophysics. Psychologists have probed human (and animal) behavior in countless experiments, and conjectured about processes that underlie memory recall, response time, attention span, interference, ability to discriminate, and so on. Relating the observations to computing with wide vectors can deepen our understanding of the underlying processes and lead to the engineering of more lifelike robots, for example.
Modeling of brains with wide vectors also opens new vistas in experimental psychology. As the theory and models improve, they will suggest experiments to confirm or counter hypothesized underlying mechanisms. This can benefit both psychology and engineering.
Language and logic are the apex of symbolic behavior and have even been considered as defining intelligence. Symbolic processing also formed the early core of AI—until we tried to build systems that learn by interacting with the world. Where logic and language failed, artificial neural nets were thought to succeed, leading to today’s deep neural nets and large language models. Impressive as they are, their energy requirements so exceed the efficiency of brains, as to rule them out as models of how brains work and make us intelligent.
@@ 9 Summary
The paper outlines a traditional (von Neumann-like) architecture for computing with wide vectors and argues for its use for autonomous learning in real time by systems such as robots. It is also meant to help us understand how humans and animals learn. It combines ideas from psychology (short-term working memory and long-term data store, chunking), neuroscience (cerebellum-like circuit for an associative memory), computer science and engineering (central processing unit, random-access memory, sequence as a linked list), and mathematics (algebra of vector operations, effective orthogonality of random vectors). Peculiar about computing with vectors is that it takes place in superposition. Variables and values are vectors of a common mathematical space, and so is a variable having a certain value. The memory for storing the vectors is also addressed by the vectors—it’s a “high-dimensional RAM.”
The subjective state of the system is defined by a single vector called the focus. It summarizes the contents of the working memory. The sensors feed into the focus, the actuators are driven from the focus, and the memory is accessed through the focus. The passage of time is represented by a sequence of vectors in the focus. When stored in the memory as a linked list, it becomes a predictive model of the world.
@@
The architecture of the long-term memory (Sparse Distributed Memory) is shallow. It has one fixed layer for activating memory locations and one variable layer for storing and recalling data in the activated locations. The fixed layer corresponds to the address-decoder circuit of a computer memory (RAM), and the variable layer corresponds to memory stacks that store the data. This architecture is suggested by the wiring of the cortex of the cerebellum, which contains over half the brain’s neurons and synapses. The variable layer can be trained in real time by the perceptron learning rule, and learning can happen in only a few rehearsals.
The vector math is carried out in the working memory (counterpart of the Central Processing Unit, CPU) with an instruction set consisting of coordinatewise addition and multiplication, and the permutation of coordinates. Similarity of vectors is based on the dot product. The operations can combine symbolic processing and statistical learning.
Interaction with the world happens through sensors and actuators. Input is made into wide vectors for the working memory/focus with sensor-specific pre-processing, and the focus is decoded for commands to actuators with motor-specific post-processing. Pre- and post-processing can be engineered, inherited or learned early in the system’s existence in multiple passes with backpropagation, and then fixed.
@@
Supplemental Material
RIACS Report 88.14 on the Organization of an Autonomous Learning System (Kanerva 1988a), referred to here as the Target Article.
Acknowledgments
This work was supported in part by CoCoSys, one of the seven centers sponsored by the Semiconductor Research Corporation (SRC) and DARPA under the Joint University Microelectronics Program 2.0 (JUMP 2.0); and in part by the U.S. National Science Foundation (NSF) under Cooperative Agreement No. 2433429, “NSF Al Research Institute on Interaction for Al Assistants (ARIA).”
Sincere thanks to professor Peter Denning for discussions that lead to writing of the paper, and to an anonymous reviewer for comments that eventually lead to its publication. Special thanks to professor Bruno Olshausen and his Redwood Center at UC Berkeley, for creating an environment favorable for scientific research and discovery.
Publisher’s Note. All claims expressed in this article are solely those of the author and do not necessarily represent those of their affiliated organizations, or those of the publisher, the editors and the reviewers. Any product that may be evaluated in this article, or claim that may be made by its manufacturer, is not guaranteed or endorsed by the publisher.